OpenStudy (nuccioreggie):

The table below shows the size of nine families selected at random from two neighborhoods in a large city: Family Size (in number of people) Neighborhood Q 2 5 4 3 2 5 3 6 5 Neighborhood S 2 3 2 3 7 2 3 3 2 Which neighborhood appears to have a bigger family size? Explain your answer using two or three sentences.

OpenStudy (michele_laino):

here you have to compute the subsequent mean values: \[\Large \begin{gathered} {m_1} = \frac{{2 + 5 + 4 + 3 + 2 + 5 + 3 + 6 + 5}}{9} = ... \hfill \\ \hfill \\ {m_2} = \frac{{2 + 3 + 2 + 3 + 7 + 2 + 3 + 3 + 2}}{9} = ... \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

please complete those computations

OpenStudy (nuccioreggie):

can you explain the answer for me please

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

it is simple, in order to answer to your question, we have to compute those mean values

OpenStudy (michele_laino):

since we are applying the statistical methods

OpenStudy (nuccioreggie):

i still havent got it yet please help @Michele_Laino

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

the bigger family size corresponds to the highest mean value

OpenStudy (michele_laino):

so, what is the right answer?

OpenStudy (nuccioreggie):

q

OpenStudy (michele_laino):

that's right!

OpenStudy (nuccioreggie):

but can you write the answer out for me explaining it please @Michele_Laino

OpenStudy (nuccioreggie):

@Michele_Laino